If Crystal gets 140 utils from eating one​ yogurt, 145 utils from eating two​ yogurts, and 148 utils from eating three​ yogurts, then​ Crystal's total utility​ is:
Answer #1

Total utility is the total satisfaction received from consuming a given total quantity of a good. Marginal utility is the change in total utility when one more unit is consumed. Crystal's TU is 140 from one yogurt, 145 from two yogurts and 148 from three yogurts.

That is, Crystal's MU from one yogurt= 140; MU from second yogurt= 145-140 = 5; MU from third yogurt= 148- 145= 3.

Thus Crystal's total utility​ is increasing and marginal utility is decreasing.